# Towards Verifiable Federated Unlearning: Framework, Challenges, and The Road Ahead

### v1

@@ Abstract
Federated unlearning (FUL) enables removing the data influence from the model trained across distributed clients, upholding the right to be forgotten as mandated by privacy regulations. FUL facilitates a value exchange where clients gain privacy-preserving control over their data contributions, while service providers leverage decentralized computing and data freshness. However, this entire proposition is undermined because clients have no reliable way to verify that their data influence has been provably removed, as current metrics and simple notifications offer insufficient assurance. We envision unlearning verification becoming a pivotal and trust-by-design part of the FUL life-cycle development, essential for highly regulated and data-sensitive services and applications like healthcare. This article introduces VeriFUL, a reference framework for verifiable FUL that formalizes verification entities, goals, approaches, and metrics. Specifically, we consolidate existing efforts and contribute new insights, concepts, and metrics to this domain. Finally, we highlight research challenges and identify potential applications and developments for verifiable FUL and VeriFUL.
@@ Introduction
    Remaining clients [^5] This article uses remaining clients and non-target clients interchangeably.: Will my contributions be intact?
It is crucial to answer these questions to establish a trust-by-design FUL system that upholds the clients’ RTBF and their right to verify the removal of data influence. Despite progress, a unified framework for verifiable FUL remains lacking. Romandini et al. in [7] surveyed unlearning algorithms and categorized metrics. Authors in [8, 9] outlined a workflow and a fine-grained taxonomy specifying who unlearns, who verifies, what is unlearned, and the key lessons and open directions. Gao et al. [10] proposed a mark-to-check protocol that allows the target client to verify the unlearning effect locally, using its hardest and most unique data samples. However, solely proposing unlearning algorithms and reporting metrics does not guarantee faithful unlearning execution. Critical gaps remain, including who conducts verification, what verification goals are across stakeholders (i.e., target clients, non-target clients, and service providers), how to implement verification, how well goals are achieved with selected verification approaches, and how to integrate these elements into a unified framework. If left unaddressed, target clients are often forced to unquestioningly trust the service providers’ and other clients’ claims about the unlearning efficacy and integrity, enabling dishonest behaviors (e.g., a service provider appears to have unlearned, passed verification, or fabricated metrics). To address this gap, we propose a unified FUL framework, called VeriFUL, which provides structured guidelines for practitioners and researchers, as well as ongoing challenges and future work in this emerging domain.
@@
This section introduces VeriFUL, specifying WHO, WHAT, HOW TO, and HOW WELL for designing a trust-by-design foundation and verifiable FUL. Figure 2 illustrates components and workflow of VeriFUL. Our focus is on the verification stage (Step 5).
@@
Clients: comprises the target and remaining clients. Target clients submit unlearning requests and retain RTBF and the right to verify that their data [^6] The data can be classified as personal or business ownership-related. influence has been effectively removed from the global model. Remaining clients verify that their contributions (e.g., updates and data impact) to the global mode remain intact. They may act as provers (e.g., proving their updates were correctly computed) and/or as verifiers, assisting checks on target-data removal.
Service providers: orchestrates learning and unlearning. When executing unlearning, the service provider must generate and provide verification artifacts such as cryptographic proofs, attestations, audit logs, and evaluation metrics, demonstrating strict adherence to the unlearning algorithm. When unlearning is performed locally by target clients, the service provider may support the secure aggregation of (unlearned) updates and assist in verifying the correctness and exclusivity of the client-side unlearning.
@@
Beyond zero-knowledge proofs, homomorphic commitments [14] and verifiable computation systems [15] can also strengthen unlearning guarantees in federated settings. Homomorphic commitments allow for a verifiable proof that the forgotten update has been correctly subtracted from the aggregation. Similarly, verifiable computation systems such as SNARKs allow the service provider to prove that it executed the prescribed unlearning algorithm correctly when recomputing the global parameters, so that clients can verify adherence to the protocol without re-running expensive computations.
@@
Auditing-based distributed ledger technology: utilizes properties of distributed ledger technologies (DLTs), such as blockchain, to create a transparent, chronological, and immutable audit trail for the entire unlearning process. Operational metadata, including unlearning requests, unlearning model updates, agreements regarding participant roles, goals, and algorithms, and verification results, are recorded as transactions on the ledger. Therefore, verifiers can verify the integrity of the unlearning process by inspecting transactions in a non-repudiable and traceable manner, thereby ensuring correctness and reversibility. However, achieving reversibility in practice requires combining DLT-based auditing with checkpointing and model-state logging, since storing full model updates on-chain is computationally expensive and economically infeasible. Instead, only cryptographic hashes of model states should be recorded on-chain [18], while the corresponding encrypted model snapshots are maintained off-chain to enable verifiable rollback when needed. Besides, the latency for achieving consensus can adversely impact the timeliness of the unlearning procedure. Deployment in regulated sectors such as healthcare also necessitates a stringent log schema design and retention policies that align with domain-specific governance frameworks, adding to the implementation complexity.
@@
Active testing: Distinct from DLT-based auditing, which provides immutable formal logs, active testing refers to empirical probing and assessment of the released unlearned model through evaluation and analysis. Instead of providing proof of correct execution like the approaches above, it reinforces trust by validating and providing the outcomes against targeted checks. For example, the verifier can test the unlearned model on chosen datasets (e.g., target client’s data), analyze differences in model updates before and after unlearning, or repurpose security and privacy attacks as audit tools (e.g., membership inference attacks-MIAs). This approach offers flexible, intuitive, and measurable results. Solely using this approach helps us to manage timeliness due to its low computational overhead, but cannot, by itself, verify correctness, as a malicious prover may fabricate the results without actually running the unlearning algorithm.
Discussion: Owning to the inherent limitations of existing verification approaches, VeriFUL demands a hybrid and goal-aware design. No single method can simultaneously optimize all core verification goals. Cryptographic proofs deliver strong guarantees for correctness and exclusivity, but harm timeliness. Active testing and auditing assess completeness and exclusivity but cannot verify process adherence. Hardware-based attestations balance correctness with timeliness but lack transparency and are vulnerable to certain attacks. Auditing-based DLTs provide a foundational mechanism for reversibility by maintaining an immutable, cryptographically-hashed log of model state transitions, but introduce significant consensus latency, impairing timeliness, and impose substantial storage overhead for recording large-scale model parameters or their hashes. Therefore, a synergistic combination is essential, leveraging the strengths of one approach to compensate for the weaknesses of another. For example, we can use cryptographic algorithms or TEEs where proofs are impractical in deployment to anchor correctness and exclusivity as a first layer; a next layer of active testing to substantiate completeness; and record minimal cryptographic hashes of model updates on a DLT to realize reversibility and enable verifiable state restoration. This multi-layered verification may sacrifice timeliness due to the computational overhead of cryptographic algorithms, the consensus latency of DLT, and the operational complexity of model and data state management, representing a necessary trade-off for achieving stronger verifiable guarantees.
@@
Verification metrics quantify how well verification goals are met under a given approach. Existing surveys summarize candidates [7, 4, 8, 9] but lack a unified standard for definitions, evaluation procedures, and acceptance criteria. Building on these efforts, VeriFUL consolidates, systematizes, and extends the metric space into a coherent taxonomy (see Step 5.4).
Completeness metrics:
  1. 1.
    Performance delta: measures the change in performance (e.g., accuracy, loss, F1 score) of the unlearned model evaluated specifically on the target data compared to the pre-unlearning model. A significant degradation in performance reflects that the model’s learned representations of target data have been removed. However, this metric alone cannot distinguish between genuine data forgetting from superficial performance suppression.
  2. 2.
    Residual-influence distance: is the discrepancy between the unlearned model and a retrained-from-scratch model trained on retained datasets, measured via (i) parameter or distribution divergence (e.g., Kullback–Leibler divergence, Wasserstein distance), representation similarity (e.g., centered kernel alignment), or weight-space distance (e.g., cosine similarity). A small distance indicates an effective approximation of full retraining.
  3. 3.
    Probe success rate: assesses the residual target data memorization by evaluating the unlearned model’s vulnerability to adversarial probes. This includes backdoor attacks, which measure the rate of embedded trigger patterns in causing targeted misclassifications. A lower rate suggests successful removal of malicious patterns. Another key metric is MIAs, wherein adversaries attempt to infer whether specific data were part of the original training set. A reduced MIA success rate on forgotten data implies enhanced unlearning efficacy and decreased privacy leakage. Additionally, influence function analysis can be used to estimate the marginal effect of the forgotten data on model predictions post-unlearning. A negligible influence score supports the efficacy of the unlearning process.
Timeliness metrics:
  1. 1.
    Latency: is the total time from request receipt to the availability of the verifiably learned model and client notification. It comprises (i) consensus time, which determines who performs unlearning and which unlearning algorithm is used; (ii) execution time, which covers running the unlearning operations on the selected entities; (iii) aggregation time, which integrates unlearned updates into a new global model; and (iv) verification time, which uses verification approaches, produces and checks proofs and logs along with verification metrics against the chosen verification goals (see Figure 2).
  2. 2.
    Throughput: is the number of unlearning requests successfully completed per unit time (e.g., per hour/day). This metric is critical for assessing system performance under high-volume, concurrent request loads.
  3. 3.
    Regulatory Deadline Adherence: is a binary or proportional score indicating compliance with legally mandated timeframes, such as 1 month for GDPR. This metric translates technical performance (i.e., latency) into a measure of regulatory risk, where 100% adherence signifies full compliance.
Correctness metrics:
  1. 1.
    Proof verification success rate (PVSR): is the proportion of valid proofs accepted by verification entities (e.g., clients, or third parties) out of the total proofs generated for a single unlearning request, whose value is in the range of 0 and 1. The PVSR of 1.0 provides deterministic evidence that the unlearning algorithm was executed correctly. Any value less than 1.0 indicates a potential compromise or error, necessitating an audit and correction.
  2. 2.
    Auditing score: is a quantitative score assigned by a third party after examining the execution logs, on-chain transactions in distributed ledger-based systems, code repositories, and learned/unlearned model checkpoints.
Exclusivity metrics:
  1. 1.
    Performance-level stability: measures the change in the performance metrics (e.g., accuracy, loss) on each remaining client’s fixed local test set using an identical evaluation protocol (e.g., data preprocessing, batch size). A value close to 0 shows that the contributions of remaining clients remain unaffected.
  2. 2.
@@
    Parameter-level stability: is the similarity (e.g., via cosine similarity) between a remaining client’s model updates before and after unlearning. Higher similarity implies intact contributions.
  3. 3.
@@
    Behaviour-level stability: measures the divergence (e.g., via Wasserstein distance) between output distributions of the unlearned global model on a client’s data before and after unlearning. Values approaching 0 indicate unchanged decision behavior.
Reversibility metrics:
  1. 1.
    Performance consistency: capture the performance difference between the restored and pre-unlearning model evaluated on a fixed test set. A negligible drift signifies high reversal fidelity.
  2. 2.
    Restoration latency: refers to the time taken to revert the unlearned model to its pre-unlearning state. Lower latency combined with performance consistency indicates an efficient reversal process.
  3. 3.
    Model state integrity: is a comparison (e.g., cryptographic hashing of model parameters) to ensure the restored model’s state is identical to the pre-unlearning, providing a deterministic guarantee of reversibility. However, this metric is often infeasible in practice due to numerical non-determinism in training operations, prohibitive costs associated with exact unlearning, or approximate unlearning.
@@
Strict enforcement of RTBF in federated settings can conflict with the need for efficient model training and deployment. When target data is highly informative or unlearning requests occur at scale, the removal and verification processes may substantially degrade the model’s generalization ability, affecting both convergence and downstream task performance. Furthermore, the additional overhead from unlearning and verification, particularly in terms of computation, communication, and coordination, delays the release of stable global models. These challenges are further magnified in resource-constrained environments where clients have limited computing, memory, and storage to support repeated and concurrent unlearning and verifiability operations.
@@
In the current landscape, a client’s intent to unlearn becomes explicitly disclosed through direct requests, which exposes the request to the service provider and other clients. While necessary for coordination and auditability, this transparency opens the door for adversarial behaviors. Particularly, the service provider may selectively omit, delay, or ignore unlearning requests, thereby undermining both the confidentiality and enforceability of the unlearning process. Moreover, using active testing-based verification (e.g., backdoor auditing and MIAs) may introduce unintended vulnerabilities [10]. Malicious verifiers (e.g., the service providers, non-target clients) could exploit these techniques to extract sensitive information, infer training set membership, or tamper with model behavior.
@@
Building on VeriFUL, we now identify four promising research directions for future investigation.
@@
While current verifiable FUL research typically assumes that unlearning is performed by the service provider and/or remaining clients [8], we envision an emerging direction that shifts the unlearning computation to the target clients. This client-centric approach alters the trust model for verification by reducing dependence on the service provider and other clients, while providing target clients greater control over unlearning processes. However, it introduces challenges in verifying that the target client’s local unlearning computation was performed correctly and exclusively on the target data. Malicious clients may exploit this process to inject poisoned updates or tamper with the global model under the guise of unlearning, posing risks to both the service provider and other clients. Moreover, FUL clients are often resource-constrained, which hinders their local unlearning and verification execution. Therefore, it is necessary to develop lightweight and scalable verification protocols for client-side unlearning.
@@
Engaging in the unlearning and verification processes imposes additional communication and computation overheads on participants. Without fair and transparent incentives, a verifiable FUL system is likely unsustainable. Frequent unlearning requests with high intensity (e.g., removal of large data volumes driven by strict privacy preferences) may cause irreversible degradation in model performance. Future research should explore incentive mechanisms that reward participants in proportion to their verifiably measured resource contributions while ensuring system-wide fairness. Particularly, designing game-theoretic-based incentive mechanisms can enable service providers and target clients to negotiate acceptable trade-offs between privacy preferences, performance degradation, and economic rewards. Such mechanisms sustain participation in training, unlearning, and verification. For instance, a transparent incentive layer atop VeriFUL framework using blockchain. Client contributions to unlearning and verification are recorded on-chain, enabling automated auditing and reward allocation via smart contracts.

### v2

@@ Abstract
Federated unlearning (FUL) enables removing the data influence from the model trained across distributed clients, upholding the right to be forgotten as mandated by privacy regulations. FUL facilitates a value exchange where clients gain privacy-preserving control over their data contributions, while service providers leverage decentralized computing and data freshness. However, this entire proposition is undermined because clients have no reliable way to verify that their data influence has been provably removed, as current metrics and simple notifications offer insufficient assurance. We envision unlearning verification becoming a pivotal and trust-by-design part of the FUL life-cycle development, essential for highly regulated and data-sensitive services and applications like healthcare. This article introduces VeriFUL, a reference framework for verifiable FUL that formalizes verification entities, goals, approaches, and metrics. Specifically, we consolidate existing efforts and contribute new insights, concepts, and metrics to this domain. Finally, we highlight research challenges and identify potential applications and developments for verifiable FUL and VeriFUL. This article aims to provide a comprehensive resource for researchers and practitioners to navigate and advance the field of verifiable FUL.
@@ Introduction
    Remaining clients [^5] This article uses remaining clients and non-target clients interchangeably.: Will my contributions remain intact?
It is crucial to answer these questions to establish a trust-by-design FUL system that upholds the clients’ RTBF and their right to verify the removal of data influence. Despite progress, a unified framework for verifiable FUL remains lacking. Romandini et al. in [15] surveyed unlearning algorithms and categorized metrics. Authors in [8, 6] outlined a workflow and a fine-grained taxonomy specifying who unlearns, who verifies, what is unlearned, and the key lessons and open directions. Gao et al. [3] proposed a mark-to-check protocol that allows the target client to verify the unlearning effect locally, using its hardest and most unique data samples. However, solely proposing unlearning algorithms and reporting metrics does not guarantee faithful unlearning execution. Critical gaps remain, including who conducts verification, what verification goals are across stakeholders (i.e., target clients, non-target clients, and service providers), how to implement verification, how well goals are achieved with selected verification approaches, and how to integrate these elements into a unified framework. If left unaddressed, target clients are often forced to unquestioningly trust the service providers’ and other clients’ claims about the unlearning efficacy and integrity, enabling dishonest behaviors (e.g., a service provider appears to have unlearned, passed verification, or fabricated metrics). To address this gap, we propose a unified FUL framework, called VeriFUL, which consolidates prior efforts [15, 8, 6, 3] on identifying who unlearns, who verifies, and how unlearning efficacy is measured, and extends them with a systematic characterization of unlearning verification goals and verification approaches, together with metrics for assessing verification efficacy and efficiency, as well as unlearning efficacy and fidelity, within a coherent framework that we then demonstrate in a healthcare case study. Furthermore, we provide ongoing challenges and future work in this emerging domain.
@@
This section introduces VeriFUL, specifying WHO, WHAT, HOW TO, and HOW WELL for designing a trust-by-design foundation and verifiable FUL. VeriFUL is a technical control layer that can be built on top of existing FUL systems, helping researchers and practitioners to implement established principles and obligations (e.g., GDPR’s RTBF, transparency, and accountability). Figure 2 illustrates components and workflow of VeriFUL with a healthcare case study. Our focus is on the verification stage (Step 5).
@@
Clients: comprise the target and remaining clients. Target clients submit unlearning requests and retain RTBF and the right to verify that their data [^6] The data can be classified as personal or business ownership-related. influence has been effectively removed from the global model. Remaining clients verify that their contributions (e.g., updates and data impact) to the global model remain intact. They may act as provers (e.g., proving their updates were correctly computed) and/or as verifiers, assisting checks on target-data removal.
Service providers: orchestrate learning and unlearning. When executing unlearning, the service provider must generate and provide verification artifacts such as cryptographic proofs, attestations, audit logs, and evaluation metrics, demonstrating strict adherence to the unlearning algorithm. When unlearning is performed locally by target clients, the service provider may support the secure aggregation of (unlearned) updates and assist in verifying the correctness and exclusivity of the client-side unlearning.
@@
Beyond ZKPs, homomorphic commitments [14] and verifiable computation systems [13] can also strengthen unlearning guarantees in federated settings. Homomorphic commitments allow for a verifiable proof that the forgotten update has been correctly subtracted from the aggregation. Similarly, verifiable computation systems such as SNARKs allow the service provider to prove that it executed the prescribed unlearning algorithm correctly when recomputing the global parameters, so that clients can verify adherence to the protocol without re-running expensive computations.
@@
Table II: Comparative Analysis of Verification Approaches in VeriFUL Framework
Approach Overhead per unlearning request Scalability w.r.t model size (M) and no. of clients (N) Security strength Use in VeriFUL
Cryptographic proofs High. Large proof size and proof generation time. Low. Degrades when M and N grow. High. Strong, formal guarantees for correctness and exclusivity. Unlearning in highly-regulated and adversarial settings with moderate M, N, and the number of unlearning requests.
Hardware-based attestation Low-moderate. Near-native execution speed and costs associated with attestation. Low-moderate. Limited by enclave memory limits; scales well with N but is unsuitable for large M [2]. Moderate. Strong guarantees for correctness (e.g., executing unlearning code and model updates/aggregation) inside the enclave, assuming the trusted hardware; vulnerable to side-channel attacks. Ensure correctness and timeliness for time-critical applications with small/moderate M; execute sensitive unlearning parts.
DLT-based auditing High. Significant storage and latency due to consensus and transaction history. Moderate. Independent of M with off-chain storage; handle large N; ledger storage and transaction confirmation latency grow with the number of unlearning events. Moderate. Strong non-repudiation and tamper-evident logging for reversibility; lacks guarantees for correctness. Regulated consortia requiring auditable history, accountability, and controlled rollback under moderate unlearning intensity.
Active testing Low. Per-verifier overhead (i.e., unlearned model evaluation and statistical/attack analysis). High. Scales well with M and N, but requires access to the final unlearned model. Low. Empirical evidence towards completeness and exclusivity, but cannot guarantee correctness itself. Periodic monitoring of the unlearning cycle, serving as a mandatory and lightweight complement to verification methods.
Distributed ledger technology-based auditing: leverages distributed ledger technologies (DLTs), such as blockchain, to create a transparent, chronological, and immutable audit trail for the unlearning workflow. In VeriFUL, a permissioned blockchain (e.g., Hyperledger Fabric v2.x [^7] https://hyperledger-fabric.readthedocs.io/en/release-2.5/), which offers higher throughput, lower latency, and fine-grained access control than public blockchains [11], is jointly operated by validator nodes (e.g., one or two anchor clients and a service provider), recording signed transactions that capture unlearning requests, role assignments, chosen learning/unlearning algorithms, key hyperparameters, model learning/unlearning checkpoints and their cryptographic commitments (e.g., hashes of model updates), as well as verification outcomes or disputes. Besides, the remaining clients that perform the unlearning act as lightweight clients. These clients submit signed transactions and can verify the audit trail without participating in the heavy consensus. Rather than storing model updates on-chain, which is technically and economically costly, the blockchain only records cryptographic hashes of model updates and key artifacts, while the corresponding encrypted artifacts remain in off-chain storage [11]. During an audit or rollback, these hashes allow verifiers to check that a recovered model state is consistent with the recorded unlearning actions. Consensus mechanisms such as Practical Byzantine Fault Tolerance [20] and its variants, which are scalable with the number of unlearning requests in terms of low latency and high throughput compared to proof of work and proof of stake, and offer deterministic finality, can be an appropriate choice for DLT-based unlearning auditing. In summary, DLT-based auditing primarily strengthens correctness and reversibility, while addressing timeliness and integrating ZKPs for mathematical correctness of unlearning execution remain important directions.
@@
Active testing: Distinct from DLT-based auditing, which provides immutable formal logs, active testing refers to empirical probing and assessment of the released unlearned model through evaluation and analysis. Instead of providing proof of correct execution like the approaches above, it reinforces trust by validating and providing the outcomes against targeted checks. For example, the verifier can test the unlearned model on chosen datasets (e.g., target client’s data), analyze differences in model updates before and after unlearning, or repurpose security and privacy attacks as audit tools (e.g., membership inference attacks-MIAs). This approach offers flexible, intuitive, and measurable results. Solely using this approach helps us to manage timeliness due to its low computational overhead, but it cannot, by itself, verify correctness, as a malicious prover may fabricate the results without actually running the unlearning algorithm.
Discussion: We analyze and summarize overhead, scalability, security, and the use in VeriFUL of verification approaches in Table II.
@@
Verification metrics quantify how well verification goals are met under a given approach. Existing surveys summarize candidates [15, 7, 8, 6] but lack a unified standard for definitions, evaluation procedures, and acceptance criteria. Building on these efforts, VeriFUL consolidates, systematizes, and extends the metric space into a coherent taxonomy (see Table III and Step 5.4 in Figure 2).
A Healthcare Case Study
To ground VeriFUL in a practical setting, we consider a healthcare consortium where multiple hospitals and cancer centers collaboratively train a global AI model for cancer treatment (e.g., the Cancer AI Alliance in the US [^8] https://www.canceralliance.ai/) without sharing raw patient data. An aggregation service, operated by the consortium or a service provider, coordinates training and performs model aggregation.
Suppose a participating hospital (i.e., the target hospital) needs to delete a subset of patient records and remove their influence on the global cancer treatment model. The target hospital submits an unlearning request to the consortium (Step 1). The consortium executes a consensus procedure to determine: (i) which entities will perform the unlearning (e.g., the aggregation service, the target hospital, and/or non-target hospitals) and (ii) which unlearning algorithm will be executed (Step 2). For example, an approximate unlearning algorithm such as FUSED [19], which identifies a few knowledge-sensitive model layers, attaches sparse trainable adapters, freezes the base model, and retrains only these adapters on the retained data, can be deployed across the aggregation service, the target hospital, and non-target hospitals to produce unlearned model updates, so that retained knowledge overwrites forgotten knowledge. Unlike standard FUL, VeriFUL ensures trust by embedding verification artifacts directly into the unlearning lifecycle. The selected entities proceed to perform the unlearning procedure (Step 3), and the aggregation service then aggregates these unlearned model updates. They generate verifiable evidence, including cryptographic commitments to model updates, training data, hyperparameters, and ZKP-based proofs of correct execution.
Therefore, VeriFUL enables a robust verification stage as shown in Step 5. Here, the participating entities agree on the set of verifiers, potentially which may include the target hospital, non-target hospitals, the aggregation service, and/or independent third-party auditors (Step 5.1). These verifiers with the target hospital in agreement then define the specific verification goals (e.g., completeness, correctness), tailored to the sensitivity of the data and consortium policies (Step 5.2). Based on these goals, they utilize the pre-generated evidence to validate the unlearning process and conduct active testing (e.g., testing the unlearned cancer treatment model performance) (Step 5.3). Finally, quantitative verification metrics (e.g., performance delta of the unlearned cancer treatment model, and PVSR) are computed and analyzed to assess how well each goal has been met, thereby providing measurable assurance to the target hospital and other stakeholders regarding the forgetting efficacy, efficiency, fidelity, and integrity of the unlearning (Step 5.4). We integrate these steps and configurations as a request-level protocol around the unlearning workflow. This makes explicit how VeriFUL can guide the design of deployable verifiable FUL pipelines, while remaining agnostic to the particular unlearning algorithms and verification primitives, thereby accommodating diverse infrastructure, application requirements, and regulatory constraints.
Table III: Taxonomy of VeriFUL Verification Metrics.
Goal Metric What it measures Typical operationalization Acceptance criteria
Completeness Performance delta Change in learned model performance before vs. after unlearning. Evaluate accuracy/loss/F1 score of unlearned model on the target dataset. Significant drop (e.g., $90\%$)
Residual-influence distance Divergence between unlearned model and a retrained-from-scratch model. Compute KL/Wasserstein distance between model distributions or parameters on retained data. Distance below predefined threshold.
Probe success rate Vulnerability to adversarial probes testing residual target data memorization. Measures backdoor ASR on triggered patterns, or MIA accuracy and influence-function scores on target data. Low ASR/accuracy/influence score.
Timeliness Latency Total time from request receipt to verified model availability. Measure end-to-end time: consensus, unlearning execution, aggregation, verification (Step 2-Step 5, see Figure 2). Meets SLA/deadline (e.g., <24<24 hours and <1<1 month for GDPR).
Throughput Number of unlearning requests completed per unit time. Aggregated logs of completed requests (e.g., per second/minute/hour/day). Throughput $\geq$ pre-defined unlearning target rate
Correctness Proof verification success rate (PVSR) Proportion of valid proofs accepted by verifiers. Generate cryptographic proofs for each unlearning step; verifiers validate. PVSR = 1.0 for all proofs. Any failure triggers audit.
Auditing score Quantitative score from third-party audit. Auditor examines logs, checkpoints, and assigns score (e.g., 0–100). Score above audit threshold (e.g., $\geq$ 90/100).
Exclusivity Performance-level stability Change in performance on remaining clients’ local test sets. Evaluate unlearned model on each remaining client’s fixed test set before/after unlearning. Performance change $\leq$ small delta (e.g., $\pm 0.5\%$ accuracy).
@@ A Healthcare Case Study
Parameter-level stability Similarity of a remaining client’s model updates before/after unlearning. May compute cosine similarity between client’s model updates. High similarity value (e.g., $\geq 0.95$ means highly intact).
@@
Behaviour-level stability Divergence in model output distributions on remaining client data before/after unlearning. May compute Wasserstein distance between output distributions. Distance value $\approx 0$ means highly intact.
Reversibility Performance consistency Performance difference between restored and pre-unlearning model. Evaluate both models on same test set. Change $\leq$ threshold (e.g., <0.5<0.5% accuracy).
Restoration latency Time to revert unlearned model to its pre-unlearning state. Measure total time from revert request to restored model availability. Latency $\leq$ latency SLO (e.g., 30 minutes).
Model state integrity Cryptographic comparison of model states pre/post-restoration. Compare hashes of model checkpoints (under deterministic training). Exact hash match (if feasible).
  • *** KL, SLA, and SLO mean Kullback-Leibler divergence, Service Level Agreement, and Service Level Objective, respectively.
  • *** Setting of pre-defined thresholds, SLAs, and SLOs are dependent on application/service/existing infrastructure/client needs.
@@
Strict enforcement of RTBF in federated settings can conflict with the need for efficient model training and deployment. When target data is highly informative or unlearning requests occur at scale, the removal and verification processes may substantially degrade the model’s generalization ability, affecting both convergence and downstream task performance. Furthermore, the additional overhead from unlearning and verification, especially using cryptographic proofs or DLTs, in terms of computation and communication, delays the release of stable global models. These challenges are further magnified in resource-constrained environments where clients have limited computing, memory, and storage to support repeated and concurrent unlearning and verifiability operations.
@@
In the current landscape, a client’s intent to unlearn becomes explicitly disclosed through direct requests, which exposes the request to the service provider and other clients. While necessary for coordination and auditability, this transparency opens the door for adversarial behaviors. Particularly, the service provider may selectively omit, delay, or ignore unlearning requests, thereby undermining both the confidentiality and enforceability of the unlearning process. Moreover, using active testing-based verification (e.g., backdoor auditing and MIAs) may introduce unintended vulnerabilities [3]. Malicious verifiers (e.g., the service providers, non-target clients) could exploit these techniques to extract sensitive information, infer training set membership, or tamper with model behavior. Therefore, designing verification mechanisms with robustness against adversarial misuse remains a pressing challenge.
@@
Building on VeriFUL, we now identify four promising research directions for future investigation, including the design of application-specific solutions, a shift towards target client-executed unlearning, the development of unlearning-driven incentive and verification mechanisms, and the formulation of predictive scaling laws for verifiable FUL. We highlight each of these directions below.
@@
While current verifiable FUL research typically assumes that unlearning is performed by the service provider and/or remaining clients [8], we envision an emerging direction that shifts the unlearning computation to the target clients. This client-centric approach alters the trust model for verification by reducing dependence on the service provider and other clients, while providing target clients greater control over unlearning processes. However, it introduces challenges in verifying that the target client’s local unlearning computation was performed correctly and exclusively on the target data. Malicious clients may exploit this process to inject poisoned updates or tamper with the global model under the guise of unlearning, posing risks to both the service provider and other clients. Moreover, FUL clients may be resource-constrained, which hinders their local unlearning and verification execution. Therefore, it is necessary to develop lightweight and scalable verification protocols for client-side unlearning.
@@
Engaging in the unlearning and verification processes imposes additional communication and computation overheads and privacy leakage risks on participants. Without fair and transparent incentives, a verifiable FUL system is likely unsustainable. Frequent unlearning requests with high intensity (e.g., removal of large data volumes driven by strict privacy preferences) may cause irreversible degradation in model performance. Future research should explore incentive mechanisms that reward participants in proportion to their verifiably measured resource contributions while ensuring system-wide fairness. Particularly, designing game-theoretic-based incentive mechanisms can enable service providers and target clients to negotiate acceptable trade-offs between privacy preferences, performance degradation, and economic rewards. Such mechanisms sustain participation in training, unlearning, and verification. For instance, a transparent incentive layer atop VeriFUL framework can be implemented using blockchain. Client contributions to unlearning and verification are recorded on-chain, enabling automated auditing and reward allocation via smart contracts.